# Multiplayer Federated Learning: Reaching Equilibrium with Less Communication

### v1

@@ 3.4 Proof outline
In this section, we provide a proof outline for Theorem 3.4. The key components of the proof are as follows: (i) a round of local SGD in PEARL-SGD behaves like a large single descent step with respect to the joint gradient operator except for local error terms caused by running multiple SGD steps locally (Lemma 3.7), and (ii) we bound these local error terms (Lemma 3.8).
Lemma 3.7.
@@
Assume (SM), and let $L_{\mathrm{max}}=\max\{L_{1},\dots,L_{n}\}$. Let $10\leq p\leq R-$ 1 be a fixed round index in PEARL-SGD and suppose $\gamma_{k}\equiv\gamma>0$ for $k=\tau p,\dots,\tau(p+1)-1$. Then for arbitrary $\alpha>0$, we have
@@
The right hand side of the bound in Lemma 3.7 involves the quantities
@@
Here we sketch the proof of Lemma 3.8 and clarify the role of Assumption (CVX). By assuming that each $f_{i}(\cdot;x_{\tau p}^{-i})$ is convex and $L_{i}$-smooth, we can prove Lemma 3.9, showing that the expectation of squared gradient norm is “almost” nonincreasing along the local SGD steps, except for some additional term due to stochasticity. Then, we rewrite each summand in (2) as
@@
and use Lemma 3.9 to bound (3).
Lemma 3.9.
Under the assumptions of Lemma 3.8, for $j=\tau p+1,\dots,\tau(p+1)$,
@@
Given (3), it is tempting to apply Jensen’s inequality to the rightmost quantity and then apply Lemma 3.9. However, this results in a bound that is looser than our Lemma 3.8. We need more careful arguments regarding the expectations, which we detail throughout Appendix B.
Proof outline for Theorem 3.4.
We combine Lemmas 3.7 and 3.8, and then apply (SCO) to eliminate the $\left\|{\varbbF}({\mathbf{x}}_{\tau p})\right\|^{2}$ terms to obtain
@@
Provided that $C\geq 0$, we can upper bound the second line of (4) by $-C\mu\left\|{\mathbf{x}}_{\tau p}-{\mathbf{x}}_{\star}\right\|^{2}$ using (QSM). Then we choose $\alpha=\gamma\tau L_{\mathrm{max}}\sqrt{\frac{\ell\mu}{3}}$ which minimizes the resulting coefficient of $\left\|{\mathbf{x}}_{\tau p}-{\mathbf{x}}_{\star}\right\|^{2}$, and rewrite it in the form $1-\gamma\tau\mu\zeta$. Finally, take expectation over ${\mathbf{x}}_{\tau p}$ in (4) and unroll the recursion. ∎
4 Numerical Experiments
In this section, we conduct experiments to assess the empirical performance of PEARL-SGD and verify our theory. We focus on three different setups: a minimax game (where $n=2$), a multiplayer game with $n=5$ players, and a distributed mobile robot control problem with $n=5$. Details of all experiments are provided in Appendix C.
4.1 Quadratic minimax game
Consider the minimax game $\min_{u\in\mathbb{R}^{d}}\max_{v\in\mathbb{R}^{d}}{\mathcal{L}}\left(u,v\right)=\frac{1}{M}\sum_{m=1}^{M}{\mathcal{L}}_{m}\left(u,v\right)$ where ${\mathcal{L}}_{m}\left(u,v\right)$ is as below (${\mathbf{A}}_{m},{\mathbf{B}}_{m},{\mathbf{C}}_{m}$ are matrices and $a_{m},c_{m}$ are vectors). In this two-player zero-sum game, we have $n=2$ with $f_{1}(x^{1};x^{2})={\mathcal{L}}(x^{1},x^{2})$ and $f_{2}(x^{2};x^{1})=-{\mathcal{L}}(x^{1},x^{2})$.
${\mathcal{L}}_{m}\left(u,v\right):=\frac{1}{2}\left\langle u,\mathbf{A}_{m}u\right\rangle+\left\langle u,\mathbf{B}_{m}v\right\rangle-\frac{1}{2}\left\langle v,\mathbf{C}_{m}v\right\rangle+\left\langle a_{m},u\right\rangle-\left\langle c_{m},v\right\rangle.$ (5)
PEARL-SGD with theoretical step-size.
Figures 2(a) and 2(b) demonstrates the performance of PEARL-SGD using the theoretical step-size $\gamma=\nicefrac{{1}}{{\left(\ell\tau+2(\tau-1)L_{\mathrm{max}}\sqrt{\kappa}\right)}}$ from Theorems 3.3 and 3.4 with $\tau\in\{1,2,4,5,8,20\}$. Figure 2(a) shows results from Deterministic PEARL-SGD, while Figure 2(b) shows the stochastic case, where we repeat each experiment 5 times and plot the mean relative error with standard deviation (shaded region). In the deterministic case, as $\gamma$ scales down with $\tau$, we observe similar linear convergence pattern for all values of $\tau$. On the other hand, in the stochastic case, we observe a clear benefit of using larger $\tau$; it reaches smaller relative error within the same number of communication rounds. This is consistent with our theory, which predicts reduced communication cost in the stochastic case.
PEARL-SGD with tuned step-size.
This experiment simulates the scenario where we do not know the precise theoretical parameters in advance. For each $\tau\in\{1,2,4,5,8,20\}$, we tune $\gamma$ by running PEARL-SGD with each $\gamma\in\{10^{-1},10^{-2},\dots,10^{-6}\}$, and plot the best relative error $\nicefrac{{\|{\mathbf{x}}_{\tau p}-{\mathbf{x}}_{\star}\|^{2}}}{{\|{\mathbf{x}}_{0}-{\mathbf{x}}_{\star}\|^{2}}}$ ($y$-axis) versus the communication round index $p$ ($x$-axis). Figure 2(c) presents results from Deterministic PEARL-SGD. We observe that performance improves as $\tau$ is increased from 1 to 5, and then degrades. Figure 2(d) presents results under stochasticity, imposed by mini-batching from the finite sum. Here we observe the lowest relative errors with large values of $\tau$, once again demonstrating the advantage of larger synchronization intervals in PEARL-SGD given stochastic gradients.
Figure 2: Performance of PEARL-SGD on quadratic minimax game (5) with different values of $\tau$. Figures 2(a) (deterministic) and 2(b) (stochastic) show the performance of PEARL-SGD with tight theoretical step-sizes, and Figures 2(c) (deterministic) and 2(d) (stochastic) show its performance with empirically tuned step-sizes.
Figure 3: Heatmap of relative errors in logarithmic scale.
Performance of PEARL-SGD for different $(\gamma,\tau)$ pairs.
Figure 3 displays the heatmap of relative errors (log-scale) after $100$ communication rounds of Deterministic PEARL-SGD on a quadratic minimax game. White and yellow regions indicate divergence/poor performance; darker regions indicate lower relative errors.
Figure 3 reveals a trend: for a fixed $\gamma$, PEARL-SGD’s performance improves as $\tau$ increases up to certain threshold, after which it declines and finally diverges. Another key observation is that the dark region of the heatmap (signifying the best performance) takes the shape of a hyperbola. This is consistent with our Theorem 3.3, showing the relationship $\gamma_{\tau}\propto\nicefrac{{1}}{{\tau}}$ where $\gamma_{\tau}$ is the optimal step-size choice given $\tau$ (providing fastest convergence).
4.2 Quadratic $n$-player game
In our second experiment, we consider an $n$-player game where the local function of the $i$-th player is given by
$f_{i}(x^{i};x^{-i}):=\frac{1}{M}\sum_{m=1}^{M}f_{i,m}(x^{i};x^{-i}),$ (6)
for $i=1,\dots,n$ (with $d_{1}=\dots=d_{n}=d$). In this setting, each $f_{i,m}$ takes the form:
$f_{i,m}(x^{i};x^{-i})=\frac{1}{2}\langle x^{i},\mathbf{A}_{i,m}x^{i}\rangle+\sum_{\begin{subarray}{c}1\leq j\leq n,j\neq i\end{subarray}}\langle x^{i},\mathbf{B}_{i,j,m}x^{j}\rangle+\langle a_{i,m},x^{i}\rangle,$
where $\mathbf{A}_{i,m},\mathbf{B}_{i,j,m}\in\mathbb{R}^{d\times d}$ and $a_{i,m}\in\mathbb{R}^{d}$ for $m=1,\dots,M$. We set the number of players to $n=5$.
Connection to game & control theory literature.
The above $n$-player game formulation has been often considered in game theory and the closely related control theory literature. It can been used to model applications including the Cournot competition in economics [2], optical networks [91], electricity markets [98], energy consumption control in smart grid [120], mobile robot control [49], etc., and has been utilized in recent works on distributed game optimization (Nash equilibrium search) [99, 119, 67, 113]. In connection with this literature, in Section 4.3, we demonstrate an experiment on a concrete robot control setup from [49].
Figure 4: Performance of PEARL-SGD on the $n$-player game defined by (6), with different $\tau$. Figures 4(a) (deterministic) and 4(b) (stochastic) show the performance of PEARL-SGD with tight theoretical step-sizes, and Figures 4(c) (deterministic) and 4(d) (stochastic) show its performance with empirically tuned step-sizes.
PEARL-SGD with theoretical step-size.
Again, we run PEARL-SGD with the theoretical step-size $\gamma=\nicefrac{{1}}{{\left(\ell\tau+2(\tau-1)L_{\mathrm{max}}\sqrt{\kappa}\right)}}$ of Theorems 3.3 and 3.4 with $\tau\in\{1,2,4,5,8,20\}$. We set the cocoercivity parameter to $\ell=\nicefrac{{L^{2}}}{{\mu}}$ following [25], where $L$ and $\mu$ are explicitly computed Lipschitz constant and strong monotonicity parameter of . Figure 4(a) displays the results from Deterministic PEARL-SGD; similarly as in Section 4.1, we observe that all values of $\tau$ produce indistinguishable performance plots. On the other hand, Figure 4(b) demonstrates that in the stochastic setting, PEARL-SGD with larger synchronization interval $\tau$ provides a clear benefit of achieving smaller relative error using the same number of communication rounds.
PEARL-SGD with tuned step-size.
In this experiment we use tuned step-size for each choice of synchronization interval $\tau\in\left\{1,2,4,5,8,20\right\}$. We use the same $\gamma$-grid $\left\{10^{-1},10^{-2},\cdots,10^{-6}\right\}$ as in Section 4.1 and proceed similarly. Figure 4(c) shows results from the deterministic setting; the choices $\tau=2$ and $\tau=20$ outperform the fully communicating case $\tau=1$ with step-size tuning. Figure 4(d) shows results from the stochastic setting, indicating that using larger values of $\tau$ could lead to higher accuracy levels.
4.3 Distributed mobile robot control
Here, we consider a distributed control problem of mobile robots from [49]. This is a multi-agent system where each robot has its own objective, depending on the positions $x^{i}\in\mathbb{R}^{d}$ (corresponding to action/strategy in our formulation of multiplayer game) of each $i$-th robot. Specifically, the objective function of the robot $i$ is:
(a) Plot of relative errors $\nicefrac{{\|{\mathbf{x}}_{\tau p}-{\mathbf{x}}_{\star}\|^{2}}}{{\|{\mathbf{x}}_{0}-{\mathbf{x}}_{\star}\|^{2}}}$
(b) Plot of local objectives $f_{i}$ ($\tau=5$)
Figure 5: Performance of PEARL-SGD on the distributed mobile robot control problem.
We implement PEARL-SGD with synchronization intervals $\tau\in\{1,2,4,5,8,20\}$ and the theoretical step-size $\gamma=\frac{1}{\ell\tau+L_{\max}(\tau-1)\sqrt{\kappa}}$. Figure 5(a) shows that with larger values of $\tau$, PEARL-SGD achieves better accuracy (in terms of distance to ${\mathbf{x}}_{\star}$) within a given number of communication rounds. This highlights the potential benefit of using local update steps in solving real-data problems formulated as multiplayer games. Figure 5(b) displays how the local objective values $f_{i}$ are reduced by PEARL-SGD, in the case $\tau=5$.
5 Conclusion
In this paper, we introduce Multiplayer Federated Learning (MpFL), a FL framework under setups where clients, strategically acting in their own interests, collaborate through a central server to train models (actions) with the goal of reaching an equilibrium. We propose the PEARL-SGD algorithm handling MpFL, and provide its tight convergence guarantees under heterogeneous setups where each player has distinct objectives and data distributions. We show that PEARL-SGD provides improved communication complexity, reducing the primary overhead in large-scale applications.
Our work offers a number of potential extensions by incorporating the ideas such as Extragradient [56, 34], asynchronous updates [17, 109], gradient compression [3], gradient tracking [90] and algorithmic correction for drifts [50, 84]. We anticipate that our initiation of the study of MpFL will lead to interesting future work including but not limited to these topics.
Acknowledgments
Sayantan Choudhury acknowledges support from the Chen Family Dissertation Fellowship. Nicolas Loizou acknowledges support from CISCO Research.
Supplementary Material
We organize the Supplementary Material as follows: Section A provides an additional survey of related work. Section B presents the proofs of theoretical results omitted from the main text. Section C provides the details of the experiments omitted from the main paper. Section D provides detailed explanation and interpretation on the theoretical assumptions made in the paper.
Appendix A Additional related work
Game Theory & Equilibrium Computation.
Multiplayer games, where multiple players each minimize their own cost function that is affected by the actions of the others, are a long-studied fundamental topic in mathematics and economics [89, 88, 102, 100, 58, 44, 76, 57, 74, 114]. More recently, there has been an increasing interest in the ML community in game-theoretic problems with motivating applications, including adversarial learning [31, 16], multi-agent reinforcement learning (MARL) [59, 66, 107], and language models [29, 46]. This incoming stream of applications has led to the development of novel analyses and insights regarding classical equilibrium-searching algorithms including gradient descent-ascent [18, 68, 117, 28, 126, 75, 60], extragradient [56, 12, 24, 62, 86, 83, 33, 34], optimistic gradient [92, 93, 94, 15, 30, 35] and consensus optimization/Hamiltonian gradient method [81, 5, 1, 75], and even the discovery of new accelerated algorithms for games [21, 121, 61, 11, 9, 122].
Heterogeneity and client drift.
One fundamental challenge for theory of Local SGD (FedAvg) is heterogeneity, i.e., varying $f_{i}$’s due to differences in local data distributions [55, 65]. Under such setup, Local SGD is prone to client drift [125, 50] where local descent trajectories head toward distinct minima (of local objectives), and convergence theories require either additional assumptions [115, 123, 39, 65] or technical analyses [51, 53] to control this drift. Some papers, based on theoretical insights, introduced or analyzed correction mechanisms for Local SGD to mitigate client drift [50, 32, 85, 84, 45, 38]. Extension of such ideas to federated minimax optimization was explored in [124].
We note that the $n$-player game setup of MpFL is also fully heterogeneous as each player has distinct (possibly even conflicting) objective functions, and consequently, we have the analogous concept of player drift. We refer the readers interested in this topic to the discussion at the end of Section 3.2.
FL frameworks with individual models.
There are several distinct contexts for FL frameworks (other than personalized FL) where each client learns an individual model. In Vertical FL [118, 72, 73] scenarios, multiple organizations hold distinct features from the common set of samples and they collaborate to train their each local model. In Federated Transfer Learning [105, 71, 27], the participating organizations similarly keep and train local models, but their datasets have heterogeneity over both sample and feature spaces with limited overlaps. Federated Multi-Task Learning [106, 77, 82] extends FL to cases where each client solves different, but related tasks.
Fictitious play.
The Fictitious Play (FP) is a classical algorithm, originally proposed by [10] to solve minimax games where each player has a finite action space and plays mixed (randomized) strategies. In FP, each player selects an action that minimizes their expected loss (best response), assuming that the other player plays the empirical (historical) strategy, which is a uniform random mixture of their previously played actions. The convergence of FP to a Nash equilibrium for minimax games was established in [95], but FP fails to converge for general $n$-player (with $n>2$) or non-zero-sum games [103, 47], except for particular cases such as all players having identical objectives [87].
While it may appear that PEARL-SGD is conceptually similar to FP (as each player performing multiple local SGD steps can be interpreted as seeking a local approximate best response to others’ strategies) the connection is opaque due to some fundamental differences. First, in PEARL-SGD, players make their updates based on only the most recent strategies of other players (not the entire history as in FP). Second, in PEARL-SGD, local SGD steps are not run until players converge to local optima—this results in player drift as we discuss at the end of Section 3.2, and is rather avoided by using step-sizes scaling down with the number of local steps. Third, in the FP setting players are assumed to have finite action spaces and mixed strategies (corresponding to points on a probability simplex), while the MpFL setting deals with continuous action spaces with pure (non-random) strategies. However, despite distinctions, as FP has been previously studied in the distributed $n$-player game setup [101], exploring the further connection between MpFL and FP could be an interesting direction.
Appendix B Omitted proofs for Per-Player Local SGD (PEARL-SGD)
B.1 Proof of Lemma 3.7
@@ Appendix B Omitted proofs for Per-Player Local SGD (PEARL-SGD)
B.2 General properties and bounds for SGD
@@
where $\mathbb{E}_{\xi}$ denotes the expectation with respect to randomness in $\xi$. This setup and the subsequent results are the abstractions of intermediate results that we need for the proofs of Lemma 3.9 and Lemma 3.8. Specifically, we will later use the results of this section with
@@
Apply Lemma B.1 recursively and use the tower rule (law of total expectation).
Lemma B.3.
@@
Lemma B.4.
Under the assumptions of Lemma B.3, we have
@@
Observe that because $\hat{x}_{t}$ as defined in Lemma B.3 is a non-random quantity and
@@
Lemma B.5.
Under the assumptions of Lemma B.3, we have
@@
where the last inequality uses Lemma B.4 (with $x_{s}$ regarded as initial point of the stochastic gradient descent). Now we apply this inequality and the induction hypothesis to (12):
@@
Next, by Jensen’s inequality and Lemma B.2,
@@
B.3 Proofs of Lemmas 3.9 and 3.8
Proof of Lemma 3.9.
@@
for the $L_{i}$-smooth convex function $f_{i}(\cdot;x_{\tau p}^{-i})$, with $x_{\tau p}^{i}$ as initial point, using the stochastic oracle $\nabla f_{i,\xi^{i}}(\cdot;x_{\tau p}^{-i})$ satisfying (BV) (unbiased estimator of $\nabla f_{i}(\cdot;x_{\tau p}^{-i})$ with variance at most $\sigma_{i}^{2}$). Therefore, we can apply Lemma B.2 with
@@
Proof of Lemma 3.8.
This is a direct consequence of Lemma B.6 with same choice of $f,\rho^{2},x_{0}$ and $L$ as in the proof of Lemma 3.9 and $x_{s}=x_{\tau p+t}^{i}$.
B.4 Remaining details in proof of Theorem 3.4
Note that the step-size condition of Lemma 3.8 is satisfied by our step-size selection, as $\gamma<\frac{2}{\ell\tau+2(\tau-1)L_{\mathrm{max}}\sqrt{\kappa}}\leq\frac{1}{L_{\mathrm{max}}(\tau-1)}$ (because $\kappa\geq 1$). Now combine Lemmas 3.7 and 3.8 to obtain
@@
Hence we can apply Theorem 3.4. Now observe that $\zeta>2-\gamma\left(\ell\tau+2(\tau-1)L_{\mathrm{max}}\sqrt{\kappa}\right)>1$, and $(1-u)^{R}\leq e^{-uR}$ for $u<1$, so
@@
Combining these with Theorem 3.4 we arrive at the desired conclusion.
B.6 Proof of Theorem 3.6
Note that we use constant step-size $\gamma_{k}\equiv\gamma_{\tau p}$ within each communication round $p$, i.e., for $\tau p\leq k\leq\tau(p+1)-1$, so we can apply the bound (18) from the proof of Theorem 3.4, provided that
@@
As $\gamma_{k}$ is constantly $\gamma_{0}=\frac{1}{\ell\tau(1+2q)}$ over rounds $p=0,\dots,p_{0}-1$, we can directly apply Theorem 3.4 with $R=p_{0}$ and similar simplification of the $\sigma^{2}$-terms as in (19) to bound
@@
Appendix C Details of Numerical Experiments
C.1 Quadratic minimax game
To ensure that the quadratic objective (5) is strongly-convex-strongly-concave and smooth, we randomly generate the symmetric matrices $\mathbf{A}_{m}$, $\mathbf{B}_{m},\mathbf{C}_{m}\in\mathbb{R}^{d\times d}$ to respectively have eigenvalues within the interval $\left[\mu_{A},L_{A}\right],\left[0,L_{B}\right]$ and $\left[\mu_{C},L_{C}\right]$, respectively, following [75]. We generate the vectors $a_{m},c_{m}\in\mathbb{R}^{d}$ to be uniformly random. In all our experiments, we set $d=10$ and $M=100$, where $M$ represents the size of dataset (number of functions in the finite sum structure). We implement PEARL-SGD with two computational nodes, one corresponding to the $x^{1}$-variable and the other to the $x^{2}$-variable.
C.2 Quadratic $n$-player game
As before, we set $d=10$ and $M=100$. The matrices $\mathbf{A}_{i,m}$ are generated randomly with their eigenvalues in the range $\left[\mu_{\mathbf{A}},L_{\mathbf{A}}\right]$ ($0<\mu_{\mathbf{A}}<L_{\mathbf{A}}$). Similarly, for $1\leq i<j\leq n$, we generate the matrices $\mathbf{B}_{i,j,m}$ randomly with their eigenvalues in $\left[0,L_{\mathbf{B}}\right]$. Notably, we set $\mathbf{B}_{j,i,m}=-\mathbf{B}_{i,j,m}^{\intercal}$ for $1\leq j<i\leq n$. With this condition, we can ensure that the $n$-player game (6) satisfies the (QSM) assumption, regardless of the values of $\mu_{\mathbf{A}},L_{\mathbf{A}}$ and $L_{\mathbf{B}}$. We show below why this is the case.

### v2

@@ C.1 Key ideas and proof outline
We first provide an outline for the proof of Theorem˜3.4. The key components of the proof are as follows: (i) we show that a round of local SGD in PEARL-SGD behaves like a large single descent step with respect to the joint gradient operator except for local error terms caused by running multiple SGD steps locally (Lemma˜C.1), and (ii) we bound these local error terms (Lemma˜C.2).
Lemma C.1.
@@
Assume (SM), and let $L_{\mathrm{max}}=\max\{L_{1},\dots,L_{n}\}$. Let $0\leq p\leq R-1$ be a fixed round index in PEARL-SGD and suppose $\gamma_{k}\equiv\gamma>0$ for $k=\tau p,\dots,\tau(p+1)-1$. Then for arbitrary $\alpha>0$, we have
@@
The right hand side of the bound in Lemma˜C.1 involves the quantities
@@
Here we sketch the proof of Lemma˜C.2 and clarify the role of Assumption (CVX). By assuming that each $f_{i}(\cdot;x_{\tau p}^{-i})$ is convex and $L_{i}$-smooth, we can prove Lemma˜C.3, showing that the expectation of squared gradient norm is “almost” nonincreasing along the local SGD steps, except for some additional term due to stochasticity. Then, we rewrite each summand in (5) as
@@
and use Lemma˜C.3 to bound (6).
Lemma C.3.
Under the assumptions of Lemma˜C.2, for $j=\tau p+1,\dots,\tau(p+1)$,
@@
Given (6), it is tempting to apply Jensen’s inequality to the rightmost quantity and then apply Lemma˜C.3. However, this results in a bound that is looser than our Lemma˜C.2. We need more careful arguments regarding the expectations, which we detail throughout Appendix˜C.
Proof outline for Theorem˜3.4.
We combine Lemmas C.1 and C.2, and then apply (SCO) to eliminate the $\left\|{\varbbF}({\mathbf{x}}_{\tau p})\right\|^{2}$ terms to obtain
@@
Provided that $C\geq 0$, we can upper bound the second line of (7) by $C\mu\left\|{\mathbf{x}}_{\tau p}-{\mathbf{x}}_{\star}\right\|^{2}$ using (QSM). Then we choose $\alpha=\gamma\tau L_{\mathrm{max}}\sqrt{\frac{\ell\mu}{3}}$ which minimizes the resulting coefficient of $\left\|{\mathbf{x}}_{\tau p}-{\mathbf{x}}_{\star}\right\|^{2}$, and rewrite it in the form $1-\gamma\tau\mu\zeta$. Finally, take expectation over ${\mathbf{x}}_{\tau p}$ in (7) and unroll the recursion. ∎
The proofs of Lemmas C.1, C.2 and C.3 and the detailed full proof of Theorem˜3.4 are presented through the following subsections.
C.2 Proof of Lemma˜C.1
@@
C.3 General properties and bounds for SGD
@@
where $\mathbb{E}_{\xi}$ denotes the expectation with respect to randomness in $\xi$. This setup and the subsequent results are the abstractions of intermediate results that we need for the proofs of Lemma˜C.3 and Lemma˜C.2. Specifically, we will later use the results of this section with
@@
Apply Lemma˜C.4 recursively and use the tower rule (law of total expectation).
Lemma C.6.
@@
Lemma C.7.
Under the assumptions of Lemma˜C.6, we have
@@
Observe that because $\hat{x}_{t}$ as defined in Lemma˜C.6 is a non-random quantity and
@@
Lemma C.8.
Under the assumptions of Lemma˜C.6, we have
@@
where the last inequality uses Lemma˜C.7 (with $x_{s}$ regarded as initial point of the stochastic gradient descent). Now we apply this inequality and the induction hypothesis to (12):
@@
Next, by Jensen’s inequality and Lemma˜C.5,
@@
C.4 Proofs of Lemmas C.3 and C.2
Proof of Lemma˜C.3.
@@
for the $L_{i}$-smooth convex function $f_{i}(\cdot;x_{\tau p}^{-i})$, with $x_{\tau p}^{i}$ as initial point, using the stochastic oracle $\nabla f_{i,\xi^{i}}(\cdot;x_{\tau p}^{-i})$ satisfying (BV) (unbiased estimator of $\nabla f_{i}(\cdot;x_{\tau p}^{-i})$ with variance at most $\sigma_{i}^{2}$). Therefore, we can apply Lemma˜C.5 with
@@
Proof of Lemma˜C.2.
This is a direct consequence of Lemma˜C.9 with same choice of $f,\rho^{2},x_{0}$ and $L$ as in the proof of Lemma˜C.3 and $x_{s}=x_{\tau p+t}^{i}$.
C.5 Remaining details in proof of Theorem˜3.4
Note that the step-size condition of Lemma˜C.2 is satisfied by our step-size selection, as $\gamma<\frac{2}{\ell\tau+2(\tau-1)L_{\mathrm{max}}\sqrt{\kappa}}\leq\frac{1}{L_{\mathrm{max}}(\tau-1)}$ (because $\kappa\geq 1$). Now combine Lemmas C.1 and C.2 to obtain
@@
Hence we can apply Theorem˜3.4. Now observe that $\zeta>2-\gamma\left(\ell\tau+2(\tau-1)L_{\mathrm{max}}\sqrt{\kappa}\right)>1$, and $(1-u)^{R}\leq e^{-uR}$ for $u<1$, so
@@
Combining these with Theorem˜3.4 we arrive at the desired conclusion.
C.7 Proof of Theorem˜3.6
Note that we use constant step-size $\gamma_{k}\equiv\gamma_{\tau p}$ within each communication round $p$, i.e., for $\tau p\leq k\leq\tau(p+1)-1$, so we can apply the bound (18) from the proof of Theorem˜3.4, provided that
@@
As $\gamma_{k}$ is constantly $\gamma_{0}=\frac{1}{\ell\tau(1+2q)}$ over rounds $p=0,\dots,p_{0}-1$, we can directly apply Theorem˜3.4 with $R=p_{0}$ and similar simplification of the $\sigma^{2}$-terms as in (19) to bound
@@
Appendix D Details of Numerical Experiments
Experiments were conducted using a personal MacBook with an Apple M3 chip and 16GB RAM.
D.1 Quadratic $n$-player game
We set $n=5$, $d=10$ and $M=100$. The matrices $\mathbf{A}_{i,m}$ are generated randomly with their eigenvalues in the range $\left[\mu_{\mathbf{A}},L_{\mathbf{A}}\right]$ ($0<\mu_{\mathbf{A}}<L_{\mathbf{A}}$). Similarly, for $1\leq i<j\leq n$, we generate the matrices $\mathbf{B}_{i,j,m}$ randomly with their eigenvalues in $\left[0,L_{\mathbf{B}}\right]$. Notably, we set $\mathbf{B}_{j,i,m}=-\mathbf{B}_{i,j,m}^{\intercal}$ for $1\leq j<i\leq n$. With this condition, we can ensure that the $n$-player game (2) satisfies the (QSM) assumption, regardless of the values of $\mu_{\mathbf{A}},L_{\mathbf{A}}$ and $L_{\mathbf{B}}$. We show below why this is the case.